solve
the proper pH of the water in swimming pools is 7.4 + or - 0.2. Write an absolute value inequality representing this range. Then solve the inequality to state the range of the pH.

Answer by Theo(13342)

let x represent the ph of the water.

|x - 7.4| <= .2

when (x - 7.4) is positve, the equation becomes (x - 7.4) <= .2 which results in x <= 7.6

when (x - 7.4) is negative, the equation becomes (x - 7.4) >= -.2 which results in x >= 7.2

you reverse the inequality and you change the sign of the expression on the right when (x - 7.4) is negative.

there's a valid reason for doing this that i won't get into unless you ask.
it involves the basic definition of absolute value.

you get x <= 7.6 and x >= 7.4.

this can be written as 7.4 <= x <= 7.6

this means that the ph is good if it reads from 7.4 to 7.6 and not good otherwise.